# Modeling extremal dependence in multivariate and spatial problems: a practical perspective

### v3

@@ 2.4.2 Bernstein polynomials modeling and Bayesian nonparametric inference
R> hyperparam <- list(mu.nbinom = 3.2, var.nbinom = 4.48)R> mm_gbp <- cbind(as.vector(mm_gbp_usd_filt), as.vector(mm_gbp_jpy_filt))R> set.seed(123)R> # Estimate dependence structure using Bayesian nonparametricR> gbp_mar <- fExtDep.np(x = mm_gbp, method = "Bayesian",+ par10 = rep(0.1, 3), par20 = rep(0.1, 3), sig10 = 0.0001, sig20 = 0.0001,+ k0 = 5, hyperparam = hyperparam, nsim = 5e+4)
Preliminary on margin 1|=====================================================================| 100%Preliminary on margin 2|=====================================================================| 100%Estimation of the extremal dependence and margins|=====================================================================| 100%
R> diagnostics(gbp_mar)The fExtDep.np routine allows for nonparametric estimation of the extremal dependence and method = "Bayesian" specifies that such a dependence is in Bernstein polynomial form and a Bayesian approach is used for the inference. The argument mar.fit = TRUE (default) allows for joint estimation of the margins and dependence while mar.prelim = TRUE (default) fits the marginal distributions using the RWMH algorithm and the fGEV routine, in order to obtain starting values for the marginal parameters. The prior distribution for $\kappa$ is set to be a negative binomial on $\kappa-3$ with mean $3.2$ and variance $4.48$, and for the point masses $p_{0}$ and $p_{1}$ uniform distributions on $[0,0.5]$ and $[a,b]$ respectively, where $a=\max\{0,(\kappa-1)p_{0}-\kappa/2+1\}$ and $b=(p_{0}+\kappa/2-1)/(\kappa-1)$. Note that both the negative binomial and the Poisson distributions are convenient choices for the prior distribution of the polynomial order $\kappa$ and the selected hyper-parameter values simply allow for a broad range of admissible values. For the prior of the point mass $p_{1}$ conditionally on $\kappa$ and $p_{0}$, the bounds of the uniform distribution are set to fulfil the conditions on the coefficients ${\bm{\eta}}_{\kappa}$. We refer to Marcon et al. (2016) for a comprehensive discussion. In this example, only the hyper-parameters need to be specified since prior.k = "nbinom" and prior.pm = "unif" by default. Lastly, a two-column object mm_gbp representing the data is created.
@@
The diagnostics function investigates the convergence of the algorithm, producing Figure 6. The top panels display the scaling parameters $\tau_{1}$ and $\tau_{2}$ from the marginal proposals and the polynomial degree $\kappa$, as function of the iterations. The bottom panels show the corresponding acceptance probability with the desired acceptance probability of $0.234$ (horizontal black solid lines). For both margins, the acceptance rate stabilizes at a level slightly below the nominated level of $0.234$ which yields decreasing values of the corresponding $\tau$ parameters. One could be tempted to increasing the number of MCMC iterations or the burn-in period but given the range of values taken, this would have little impact. Overall, Figure 6 suggests a burn-in period of $30,000$ iterations and we now run the following commands
R> # Compute summary statistics of the MCMC outputR> gbp_mar_sum <- summary(object = gbp_mar, burn = 30000, plot = TRUE)R> mm_gbp_rg <- apply(mm_gbp, 2, quantile, c(0.9, 0.995))R> y_gbp_usd <- seq(from = mm_gbp_rg[1, 1], to = mm_gbp_rg[2, 1], length = 20)R> y_gbp_jpy <- seq(from = mm_gbp_rg[1, 2], to = mm_gbp_rg[2, 2], length = 20)R> y <- as.matrix(expand.grid(y_gbp_usd, y_gbp_jpy, KEEP.OUT.ATTRS = FALSE))R> # Compute and plot return levelsR> ret_marg <- returns(x = gbp_mar, summary.mcmc = gbp_mar_sum, y = y,+ plot = TRUE, data = mm_gbp, xlab = "GBP/USD exchange rate",+ ylab = "GBP/JPY exchange rate")The summary routine applied to objects of class ExtDep_npBayes computes summary statistics of the MCMC output, including posterior sample, mean and $95\%$ credibility intervals (the latter can be modified using argument cred) for the angular density and Pickands dependence function. Posterior mean and credibility intervals are also computed for all parameters. Setting plot=TRUE displays the posterior mean and credibility intervals for both angular density and Pickands dependence function, as well as the prior and posterior distribution for the point mass $p_{0}$ and the polynomial degree $\kappa$. The returns routine inputs objects of class ExtDep_npBayes and outputs of the summary function (via the arguments x and summary.mcmc) to compute exceeding probabilities as defined in (2.17), for extreme values specified by y. The argument plot calls the plot routine on the object x to visualize such probabilities as long as y defines a square grid and data adds the relevant datapoints. Usage of the summary and returns functions is provided below with graphical outputs presented in Figure 7.
@@
As mentioned at the beginning of this section, computing small conditional probabilities such as ${\mathbb{P}}(\textrm{GBP/USD}>q_{1}\mid\textrm{GBP/JPY}>q_{2})$} or ${\mathbb{P}}(\textrm{GBP/JPY}>q_{2}\mid\textrm{GBP/USD}>q_{1})$} can be of interest. To proceed, take $(q_{1},q_{2})\approx(0.0069,0.0102)$ which corresponds to the observed $99\%$ quantiles. The joint probability of exceedance is computed (through returns) using the approximation ${\mathbb{P}}(\textrm{GBP/USD}>q_{1},\textrm{GBP/JPY}>q_{2})\approx{\mathbb{P}}(Z_{1}>z_{1}^{*},Z_{2}>z_{2}^{*})\approx 0.0098$, where $z_{1}^{*}$ and $z_{2}^{*}$ are the transformation of $q_{1}$ and $q_{2}$ to unit Fréchet scale. Using the pGEV function to evaluate the marginal probability of exceedance, we obtain ${\mathbb{P}}(\textrm{GBP/USD}>q_{1}\mid\textrm{GBP/JPY}>q_{2})\approx 0.3078$ and ${\mathbb{P}}(\textrm{GBP/JPY}>q_{2}\mid\textrm{GBP/USD}>q_{1})\approx 0.3804$. The computations are presented in the code below.
R> qs <- apply(mm_gbp, 2, quantile, c(0.99))R> # Compute joint probability of exceedanceR> jointP <- returns(out = gbp_mar, summary.mcmc = gbp_mar_sum,+ y = matrix(qs, ncol = 2))R> # Compute conditional probabilities of exceedanceR> par1 <- gbp_mar_sum\$mar1.meanR> par2 <- gbp_mar_sum\$mar2.meanR> jointP / pGEV(qs[1], loc = par1[1], scale = par1[2], shape = par1[3],+ lower.tail = FALSE)R> jointP / pGEV(qs[2], loc = par2[1], scale = par2[2], shape = par2[3],+ lower.tail = FALSE)The model’s goodness of fit provided can be verified graphically at the marginal and dependence level. First, marginal density histograms can be plotted together with GEV density curves evaluated at the posterior samples of the marginal parameters. Second, after standardizing the data to unit Fréchet marginals, a density histogram of the angular components with the largest radial component can be used as an empirical estimate of the angular density, to which the model estimates (posterior mean, $0.05$ and $0.95$ quantile) can be added. Figure 14, given in Appendix A.1, provides graphical evidence that the proposed model does provide a good fit to the data.
@@ 2.4.3 Simulation of extreme values
For simplicity, most of the simulation methods assume that the extremal dependence complies with some parameter model, while few attempt to consider a nonparametric approach. Marcon et al. (2017) proposed a flexible procedure for sampling from bivariate extremes with a semiparametric dependence structure, which is summarized as follows. For every small probabilities $p_{1}$ and $p_{2}$ such that $L(p_{1},p_{2})\in[0,1]$ and $R(p_{1},p_{2})\in[0,1]$, the tail probabilities in (2.9) and (2.10) can be approximated by ${\mathbb{P}}\{X_{1}>Q_{1}(p_{1})\,\text{or}\,X_{2}>Q_{2}(p_{2})\}\approx L(p_{1},p_{2})$ and ${\mathbb{P}}\{X_{1}>Q_{1}(p_{1})\,\text{and}\,X_{2}>Q_{2}(p_{2})\}\approx R(p_{1},p_{2})$. For such probabilities $p_{1}$ and $p_{2}$, we have that $L(p_{1},p_{2})=2{\mathbb{E}}\{\max(p_{1}W,p_{2}(1-W)\}$ and $R(p_{1},p_{2})=2{\mathbb{E}}\{\min(p_{1}W,p_{2}(1-W)\}$, where $W$ is a random variable on $[0,1]$ distributed according to the angular distribution $H$. Let $(Z_{1},Z_{2})=R(W,1-W)$, where $R$ is a unit-Pareto random variable, then ${\mathbb{P}}(Z_{1}>z_{1}\,\text{or}\,Z_{2}>z_{2})=2{\mathbb{E}}\{\max(W/z_{1},(1-W)/z_{2})\}$ and ${\mathbb{P}}(Z_{1}>z_{1}\,\text{and}\,Z_{2}>z_{2})=2{\mathbb{E}}\{\min(W/z_{1},(1-W)/z_{2})\}$. Marcon et al. (2017) proposed to model $H$ through Bernstein polynomials and demonstrated that $H$ can be written as a finite mixture of beta distributions with weights defined by a suitable transformation of the polynomial coefficients. This led to a simple algorithm to sample from $H$ (see Algorithm 1) and an algorithm for sampling observations from the tail of a bivariate distribution (see Algorithm 3), which can be used to approximate the corresponding tail probabilities. Briefly, from (2.6) we have that for a sufficiently large $n$ and a sufficiently small $p_{j}=p_{n,j}$ (with $p_{j}\to 0$ as $n\to\infty$), $Q_{j}(p_{j})\approx\mu_{j}+\gamma_{j}\{(np_{j})^{-\gamma_{j}}-1\}/\sigma_{j}$, with $j=1,2$. Let $u_{1}$ and $u_{2}$ be two high thresholds such that, for every $p_{1}$ and $p_{2}$ for which the above approximations hold, and in addition, $L(p_{1},p_{2})\in[0,1]$ and $R(p_{1},p_{2})\in[0,1]$, we have $Q_{j}(p_{j})>u_{j}$, $j=1,2$. For the failure regions $\mathcal{A}_{\bm{u}}=\left\{(v_{1},v_{2}):v_{1}>u_{1}\textrm{ or }v_{2}>u_{2}\right\}$ and $\mathcal{B}_{\bm{u}}=\left\{(v_{1},v_{2}):v_{1}>u_{1}\textrm{ and }v_{2}>u_{2}\right\}$ we have ${\mathbb{P}}({\bm{X}}\in\mathcal{A}_{\bm{u}})\approx{\mathbb{P}}(Z_{1}>u^{\star}_{1}\,\text{or}\,Z_{2}>u^{\star}_{2})$ and ${\mathbb{P}}({\bm{X}}\in\mathcal{B}_{\bm{u}})\approx{\mathbb{P}}(Z_{1}>u^{\star}_{1}\,\text{and}\,Z_{2}>u^{\star}_{2})$, where $u^{\star}_{j}=(1+\gamma_{j}(u_{j}-\mu_{j})/\sigma_{j})^{1/\gamma_{j}}$, for $j=1,2$. By simulating a large sample of angular components $(w_{1},\ldots,w_{N})$ from $H$ and radial components $(r_{1},\ldots,r_{N})$ from a unit-Pareto distribution, we compute ${\bm{z}}_{i}=2(r_{i}w_{i},r_{i}(1-w_{i}))$, for $i=1,\ldots,N$ and estimate of the probability of falling in $\mathcal{A}_{\bm{u}}$ and $\mathcal{B}_{\bm{u}}$ by
@@
The angular distribution is then estimated using the approximate likelihood approach (see Beranger and Padoan, 2015). Observations with radius component greater than their $90\%$ quantile are selected and the routine fExtDep.np is then called with arguments method = "Frequentist" and type = "maxima" to maximize the likelihood of a polynomial angular distribution in Bernstein form through the non-linear optimization routine nloptr from the nloptr package (Ypma and Johnson, 2022) subject to constraints established in Marcon et al. (2017). Empirical studies (Marcon et al., 2017) suggest that the polynomial degree $\kappa=10$ is enough to represent flexible dependence structures. When type = "rawdata, data are extracted using a threshold on the radial component set by the argument u and the likelihood for a sample maxima written as function of the Pickands dependence function in Bernstein form is optimized. In addition, when mar.fit = TRUE then marginal empirical transformation to unit Fréchet of the data is applied. The plot routine with type = "summary" displays graphical summaries (Pickands dependence function and angular density) of the estimated dependence structure (see left and middle panels of Figure 8). A moderate level of dependence is observed as well as point masses at the vertices. Given that the middle panel of Figure 8 combines empirical information with a summary of the fitted model, it can be leveraged to highlight the good fit of the model. The routine rExtDep with arguments model = "semi.bvevd" and angular = TRUE generates pseudo-angles according to Algorithm 1 of Marcon et al. (2017). The right panel of Figure 8 presents a histogram of $1,000$ randomly generated angles and point-masses (black triangles), the red line and dots represent the estimated dependence structure. The corresponding code is

### v5

@@ 2.4.2 Bernstein polynomials modeling and Bayesian nonparametric inference
R> hyperparam <- list(mu.nbinom = 3.2, var.nbinom = 4.48)R> mm_gbp <- cbind(as.vector(mm_gbp_usd_filt), as.vector(mm_gbp_jpy_filt)) * 100R> set.seed(123)R> # Estimate dependence structure using Bayesian nonparametricR> par0 <- c(-0.1, 0.1, -0.1)R> gbp_mar <- fExtDep.np(x = mm_gbp, method = "Bayesian", par10 = par0,+ par20 = par0, sig10 = 0.1, sig20 = 0.1, k0 = 5,+ hyperparam = hyperparam, nsim = 5e+4)
R> diagnostics(gbp_mar)Note that the data has been rescaled by a factor of $100$ to avoid numerical instabilities due to the small scale of the data. The fExtDep.np routine allows for nonparametric estimation of the extremal dependence and method = "Bayesian" specifies that such a dependence is in Bernstein polynomial form and a Bayesian approach is used for the inference. The argument mar.fit = TRUE (default) allows for joint estimation of the margins and dependence while mar.prelim = TRUE (default) fits the marginal distributions using the RWMH algorithm and the fGEV routine, in order to obtain starting values for the marginal parameters. The prior distribution for $\kappa$ is set to be a negative binomial on $\kappa-3$ with mean $3.2$ and variance $4.48$, and for the point masses $p_{0}$ and $p_{1}$ uniform distributions on $[0,0.5]$ and $[a,b]$ respectively, where $a=\max\{0,(\kappa-1)p_{0}-\kappa/2+1\}$ and $b=(p_{0}+\kappa/2-1)/(\kappa-1)$. Note that both the negative binomial and the Poisson distributions are convenient choices for the prior distribution of the polynomial order $\kappa$ and the selected hyper-parameter values simply allow for a broad range of admissible values. For the prior of the point mass $p_{1}$ conditionally on $\kappa$ and $p_{0}$, the bounds of the uniform distribution are set to fulfill the conditions on the coefficients ${\bm{\eta}}_{\kappa}$. We refer to Marcon et al. (2016) for a comprehensive discussion. In this example, only the hyper-parameters need to be specified since prior.k = "nbinom" and prior.pm = "unif" by default. Lastly, a two-column object mm_gbp representing the data is created.
@@
The diagnostics function investigates the convergence of the algorithm, producing Figure 6. The top panels display the scaling parameters $\tau_{1}$ and $\tau_{2}$ from the marginal proposals and the polynomial degree $\kappa$, as function of the iterations. The bottom panels show the corresponding acceptance probability with the desired acceptance probability of $0.234$ (horizontal black solid lines). For both margins, the acceptance rate stabilizes at a level slightly below the nominated level of $0.234$ which yields decreasing values of the corresponding $\tau$ parameters. One could be tempted to increasing the number of MCMC iterations or the burn-in period but given the range of values taken, this would have little impact. In addition, the first two rows of the Figure 6 indicate that acceptance probability for the marginal estimates is slightly below the nominated levels. Actual convergence of the marginal chains can be confirmed using traceplots on the columns of the objects gbp_mar\$mar1 and gbp_mar\$mar2. Overall, Figure 6 suggests a burn-in period of $30,000$ iterations and we now run the following commands
R> # Compute summary statistics of the MCMC outputR> gbp_mar_sum <- summary(object = gbp_mar, burn = 30000, plot = TRUE)R> mm_gbp_rg <- apply(mm_gbp, 2, quantile, c(0.9, 0.995))R> y_gbp_usd <- seq(from = mm_gbp_rg[1, 1], to = mm_gbp_rg[2, 1], length = 20)R> y_gbp_jpy <- seq(from = mm_gbp_rg[1, 2], to = mm_gbp_rg[2, 2], length = 20)R> y <- as.matrix(expand.grid(y_gbp_usd, y_gbp_jpy, KEEP.OUT.ATTRS = FALSE))R> # Compute and plot return levelsR> ret_marg <- returns(x = gbp_mar, summary.mcmc = gbp_mar_sum, y = y,+ plot = TRUE, data = mm_gbp, xlab = "GBP/USD exchange rate (x100)",+ ylab = "GBP/JPY exchange rate (x100)")The summary routine applied to objects of class ExtDep_npBayes computes summary statistics of the MCMC output, including posterior sample, mean and $95\%$ credibility intervals (the latter can be modified using argument cred) for the angular density and Pickands dependence function. Posterior mean and credibility intervals are also computed for all parameters. Setting plot=TRUE displays the posterior mean and credibility intervals for both angular density and Pickands dependence function, as well as the prior and posterior distribution for the point mass $p_{0}$ and the polynomial degree $\kappa$. The returns routine inputs objects of class ExtDep_npBayes and outputs of the summary function (via the arguments x and summary.mcmc) to compute exceeding probabilities as defined in (2.17), for extreme values specified by y. The argument plot calls the plot routine on the object x to visualize such probabilities as long as y defines a square grid and data adds the relevant datapoints. Usage of the summary and returns functions is provided below with graphical outputs presented in Figure 7.
@@
As mentioned at the beginning of this section, computing small conditional probabilities such as ${\mathbb{P}}(\textrm{GBP/USD}>q_{1}\mid\textrm{GBP/JPY}>q_{2})$} or ${\mathbb{P}}(\textrm{GBP/JPY}>q_{2}\mid\textrm{GBP/USD}>q_{1})$} can be of interest. We proceed by taking $(q_{1},q_{2})\approx(0.0069,0.0102)$ which corresponds to the observed marginal $99\%$ quantiles. The joint probability of exceedance is computed (through returns) using the approximation ${\mathbb{P}}(\textrm{GBP/USD}>q_{1},\textrm{GBP/JPY}>q_{2})\approx{\mathbb{P}}(Z_{1}>z_{1}^{*},Z_{2}>z_{2}^{*})\approx 0.0044$, where $z_{1}^{*}$ and $z_{2}^{*}$ are the transformation of $q_{1}$ and $q_{2}$ to unit Fréchet scale. Using the pGEV function to evaluate the marginal probability of exceedance, we obtain ${\mathbb{P}}(\textrm{GBP/USD}>q_{1}\mid\textrm{GBP/JPY}>q_{2})\approx 0.3169$ and ${\mathbb{P}}(\textrm{GBP/JPY}>q_{2}\mid\textrm{GBP/USD}>q_{1})\approx 0.2767$. The computations are presented in the code below.
R> qs <- apply(mm_gbp, 2, quantile, c(0.99))R> # Compute joint probability of exceedanceR> jointP <- returns(x = gbp_mar, summary.mcmc = gbp_mar_sum,+ y = matrix(qs, ncol = 2))R> # Compute conditional probabilities of exceedanceR> par1 <- gbp_mar_sum\$mar1.meanR> par2 <- gbp_mar_sum\$mar2.meanR> jointP / pGEV(qs[1], loc = par1[1], scale = par1[2], shape = par1[3],+ lower.tail = FALSE)R> jointP / pGEV(qs[2], loc = par2[1], scale = par2[2], shape = par2[3],+ lower.tail = FALSE)The model’s goodness of fit provided can be verified graphically at the marginal and dependence level. First, marginal density histograms can be plotted together with GEV density curves evaluated at the posterior samples of the marginal parameters. Second, after standardizing the data to unit Fréchet marginals, a density histogram of the angular components with the largest radial component can be used as an empirical estimate of the angular density, to which the model estimates (posterior mean, $0.05$ and $0.95$ quantile) can be added. Figure 14, given in Appendix A.1, provides graphical evidence that the proposed model does provide a good fit to the data.
@@ 2.4.3 Simulation of extreme values
For simplicity, most of the simulation methods assume that the extremal dependence complies with some parameter model, while few attempt to consider a nonparametric approach. Marcon et al. (2017) proposed a flexible procedure for sampling from bivariate extremes with a semiparametric dependence structure, which is summarized as follows. For every small probabilities $p_{1}$ and $p_{2}$ such that $L(p_{1},p_{2})\in[0,1]$ and $R(p_{1},p_{2})\in[0,1]$, the tail probabilities in (2.9) and (2.10) can be approximated by ${\mathbb{P}}\{X_{1}>Q_{1}(p_{1})\,\text{or}\,X_{2}>Q_{2}(p_{2})\}\approx L(p_{1},p_{2})$ and ${\mathbb{P}}\{X_{1}>Q_{1}(p_{1})\,\text{and}\,X_{2}>Q_{2}(p_{2})\}\approx R(p_{1},p_{2})$. For such probabilities $p_{1}$ and $p_{2}$, we have that $L(p_{1},p_{2})=2{\mathbb{E}}\{\max(p_{1}W,p_{2}(1-W)\}$ and $R(p_{1},p_{2})=2{\mathbb{E}}\{\min(p_{1}W,p_{2}(1-W)\}$, where $W$ is a random variable on $[0,1]$ distributed according to the angular distribution $H$. Let $(Z_{1},Z_{2})=R(W,1-W)$, where $R$ is a unit-Pareto random variable, then ${\mathbb{P}}(Z_{1}>z_{1}\,\text{or}\,Z_{2}>z_{2})=2{\mathbb{E}}\{\max(W/z_{1},(1-W)/z_{2})\}$ and ${\mathbb{P}}(Z_{1}>z_{1}\,\text{and}\,Z_{2}>z_{2})=2{\mathbb{E}}\{\min(W/z_{1},(1-W)/z_{2})\}$. Marcon et al. (2017) proposed to model $H$ through Bernstein polynomials and demonstrated that $H$ can be written as a finite mixture of beta distributions with weights defined by a suitable transformation of the polynomial coefficients. This leads to a simple algorithm to sample from $H$ (see Algorithm 1) and an algorithm for sampling observations from the tail of a bivariate distribution (see Algorithm 3), which can be used to approximate the corresponding tail probabilities. Briefly, from (2.6) we have that, for sufficiently large $n$ and sufficiently small $p_{j}=p_{n,j}$ (with $p_{j}\to 0$ as $n\to\infty$), $Q_{j}(p_{j})\approx\mu_{j}+\gamma_{j}\{(np_{j})^{-\gamma_{j}}-1\}/\sigma_{j}$, for $j=1,2$. Let $u_{1}$ and $u_{2}$ be two high thresholds such that, for every $p_{1}$ and $p_{2}$ for which the above approximations hold, and for which in addition, $L(p_{1},p_{2})\in[0,1]$ and $R(p_{1},p_{2})\in[0,1]$, we have $Q_{j}(p_{j})>u_{j}$, $j=1,2$. For the failure regions $\mathcal{A}_{\bm{u}}=\left\{(v_{1},v_{2}):v_{1}>u_{1}\textrm{ or }v_{2}>u_{2}\right\}$ and $\mathcal{B}_{\bm{u}}=\left\{(v_{1},v_{2}):v_{1}>u_{1}\textrm{ and }v_{2}>u_{2}\right\}$ we have ${\mathbb{P}}({\bm{X}}\in\mathcal{A}_{\bm{u}})\approx{\mathbb{P}}(Z_{1}>u^{\star}_{1}\,\text{or}\,Z_{2}>u^{\star}_{2})$ and ${\mathbb{P}}({\bm{X}}\in\mathcal{B}_{\bm{u}})\approx{\mathbb{P}}(Z_{1}>u^{\star}_{1}\,\text{and}\,Z_{2}>u^{\star}_{2})$, where $u^{\star}_{j}=(1+\gamma_{j}(u_{j}-\mu_{j})/\sigma_{j})^{1/\gamma_{j}}$, for $j=1,2$. By simulating a large sample of angular components $(w_{1},\ldots,w_{N})$ from $H$ and radial components $(r_{1},\ldots,r_{N})$ from a unit-Pareto distribution, we compute ${\bm{z}}_{i}=2(r_{i}w_{i},r_{i}(1-w_{i}))$, for $i=1,\ldots,N$ and estimate of the probability of falling in $\mathcal{A}_{\bm{u}}$ and $\mathcal{B}_{\bm{u}}$ by
@@
The angular distribution is then estimated using the approximate likelihood approach (see Beranger and Padoan, 2015). Observations with radius component greater than their $90\%$ quantile are selected and the routine fExtDep.np is then called with arguments method = "Frequentist" and type = "maxima" to maximize the likelihood of a polynomial angular distribution in Bernstein form through the non-linear optimization routine nloptr from the nloptr package (Ypma and Johnson, 2025) subject to constraints established in Marcon et al. (2017). Empirical studies (Marcon et al., 2017) suggest that the polynomial degree $\kappa=10$ is enough to represent flexible dependence structures. When type = "rawdata, data are extracted using a threshold on the radial component set by the argument u and the likelihood for a sample maxima written as function of the Pickands dependence function in Bernstein form is optimized. In addition, when mar.fit = TRUE then marginal empirical transformation to unit Fréchet of the data is applied. The plot routine with type = "summary" displays graphical summaries (Pickands dependence function and angular density) of the estimated dependence structure (see left and middle panels of Figure 8). A moderate level of dependence is observed as well as point masses at the vertices. Given that the middle panel of Figure 8 combines empirical information with a summary of the fitted model, it can be leveraged to highlight the good fit of the model. The routine rExtDep with arguments model = "semi.bvevd" and angular = TRUE generates pseudo-angles according to Algorithm 1 of Marcon et al. (2017). The right panel of Figure 8 presents a histogram of $1,000$ randomly generated angles and point-masses (black triangles), the red line and dots represent the estimated dependence structure. The corresponding code is